OpenStudy (anonymous):

What is the exact value of Sin 5pi/2. I need to know some steps on how to get the answer. the answers are A.0 B.1 C. Undefined. D. 1/2. please help

OpenStudy (anonymous):

Uhm, you can try using a calculator.

OpenStudy (anonymous):

Thanks, but the question is requesting for an exact answer that can't be found with a calculator.

OpenStudy (asnaseer):

\(\frac{5\pi}{2}\) is an exact value. Unless the question is stating that this is a radian measure and asking for the equivalent value in degrees? In which case use the fact that \(\frac{\pi}{2}=90^\circ\) so \(\frac{5\pi}{2}=5*90^\circ=450^\circ\)

OpenStudy (anonymous):

oops i forgot to put \[\sin 5\pi/2\]

OpenStudy (asnaseer):

Use the fact that the sine curve repeats itself every \(360^\circ\) (i.e. every \(2\pi\) radians)

OpenStudy (anonymous):

Thanks i will try that